# Agentic Meta-Orchestrator for Multi-task Copilots

### v1

@@ I Introduction
Microsoft copilots are designed to handle multiple work tasks by coordinating large language models (LLMs) [1, 2, 3, 4], Microsoft 365 productivity apps, and more. LLMs are impressive at handling general tasks, such as drafting documents, question answering, and summarization. Due to data compliance and boundaries, fine-tuning of customer data with web-based LLMs, such as ChatGPT, can be prohibited. Therefore, keeping the knowledge of an LLM up-to-date can be challenging.
@@
To provide a customized in-domain assistant for Microsoft customers, language agents can neatly extend the generalizability of LLMs in private domains [5, 6, 7, 8] and connect with applications, knowledge graphs, databases, other copilots, etc. The explosion of adapting various agents in particular tasks or domains poses challenges for 1) an effective orchestrator that decomposes tasks from user prompts and assigns them to growing agents, 2) efficient deployment of the foundation model and agents, and 3) inference planning among agents, e.g., which agents should be used next, and what messages can be shared across agents.
Regarding the first challenge, we propose a trained multi-level rating learning-to-rank model (with graded relevance) to orchestrate the prompts to the right agents, which is generally taken as a hierarchical text classification problem [9, 10, 11]. The agents can be “ask for price”, “compare products”, “contact human customer support”, etc. as in our M365 E-Commerce Copilot, and can be ”encoding character issue”, ”new character issue” as in the code compliance copilot. An orchestrator can be at the copilot level, e.g., it selects the right plugins  [^1] https://learn.microsoft.com/en-us/microsoft-365-copilot/extensibility/orchestrator with agents, and it can select the right models based on the business intents of user prompts. As the number of agents continues to grow, the use of a text classification model [12, 13] can have difficulties in scalability, and the cosine similarity method can struggle when the descriptions of agents have overlaps or when user prompts are ambiguous.
Regarding the second challenge, we propose a memory-efficient inference framework, LoRA[14] arms, where each arm handles one task. Hosting multiple LLMs requires substantial memory[15, 16]. Inspired by the LoRA framework, we share the base LLM model among tasks and train LoRA weights for individual tasks.
Regarding the third challenge, we propose a novel meta-learning[17] decision tree model that trains the best inference model combinations and orderings for different prompts. The well-known successful framework, retrieval-augmented generation (RAG), is a simplified two-step inference planning based on human heuristics that can be taken as a web search agent and LLM question-answering agent. When a complex copilot service is composed of multiple agents and combinations of skills/models, how we effectively choose an inference strategy for individual user prompts requires a meta-learning approach. Current planning strategies, such as AutoGen[18, 19, 20, 21, 22], CrewAI, and TaskWeaver[23], rely on agentic workflows and task-specific prompts, which are generally based on cognitive architectures[24, 25, 26, 27].
@@ II-A Task Distribution and Semantic Learning-to-rank
LLMs have been attempted in few-shot text classification tasks[28, 29] and few-shot learning-to-rank models[30, 31, 32, 33]
Previous studies have shown that learning-to-rank can be used explicitly and implicitly in various language tasks [34, 35]. However, they were limited to binary ratings. We apply semantic learning to rank to multi-level rating scenarios. We utilize the natural description of classification labels as the learning-to-rank “candidate documents”, different from the 0 to $n$ classes using one-hot encoding in softmax.
@@ II-B Multi-task Learning
Multi-task learners generally share a common module [36, 11]. Previous studies, such as MT-DNN[37] and MulT[38] have focused on training multiple tasks at the same time by sampling one task at a time using a neural network with shared layers and task-specific layers. This method can suffer performance drops in subsets of tasks, and it is not easy to efficiently update models based on individual task requirements. The authors of [39] proposed projected attention layers (PALs) in parallel to each layer for BERT [13], which learns task-specific information in the additional parameters. The idea is similar to LoRA, which uses low-rank matrices to reduce the number of parameters for attention in task-specific layers. However, when we update the training data for one task, the performance of other tasks also changes, which is not good for updating the models, especially considering in many cases in industry we want to update the model for one task without affecting the performance for other tasks. We show that we achieve impressive individual and end-to-end task performance leveraging LoRA fine-tuning and inference memory optimization in Sections III-B - III-C and IV-C.
@@ II-C Agentic Workflow
Unlike chain of thoughts[40], tree of thoughts [41], or tree search [42, 43] that address LLM inference planning, our meta-learning planning strategy is a trained model of models/agents.
A well-known cognitive heuristics RAG [44] plans a knowledge retrieval then LLM generation. Although great reinforcement learning approaches, for example, React [45] and Reflection [46], and Self-critic [47] have been attempted for complex systems, how we quickly update the planning when new agents are added remains an open problem. Our proposed AMO novelly solves this problem of continually adding new agents.
Another good cognitive heuristic is ensemble modeling. Each record is processed by the same models using bagging, stacking, etc. Our AMO, in contrast, teaches the methodology of training a model [48, 49, 50, 29] and is Conversable; agents can share information with each other.
@@ III-A Agentic Orchestrator of Top-k Selections
Therefore, we design an orchestrator that can seamlessly take new agents on the basis of their agent/class labels. Figure 2 demonstrates that we convert agent/class descriptions to semantic embeddings offline and use them to train a listwise multi-level learning-to-rank model [51, 52, 53]. The model targets higher relatedness prediction scores for agents that are ranked higher than others. In particular, when there is only one correct item in a ranking list or classification group, such as ‘[0, 1, 0, 0]’, the gradients of ListMLE-style loss functions and Softmax are equivalent. However, the former is more applicable to multi-level ratings or hierarchical classes. The user prompt and the agents use the same embedding model, which is generic. We were able to achieve stable performance using BERT and Sentence Transformers as the foundation models, even more stable than using softmax classification trained for all classes.
Each batch contains the trainable embedding of $q$ a user prompt, fixed embeddings of $D^{q}=\{d\}$ the set of candidate agent descriptions associated with $q$, which can differ in batches, and $R(D^{q})=\{r(d)\}$ the relatedness rating for $D^{q}$, ranging over all real numbers (showing that the previous study in [53] can be generalized). We maximize the likelihood of selecting a document or class $d$ from $c$ = $d\cup\tilde{s}$:
@@
The rating of every document in $\tilde{s}$ is less than r(d)r(d)[53]. We finalize our general learning to rank loss function uRank loss by adding a weight factor of $2^{r(d)}-1$. We list and answer the following core questions accordingly.
@@
LoRA-Arms are trained eastern standardduring fine-tuning and used simultaneously during inference.
@@ III-B LoRA Arms
Slightly different from LoRA memory optimization that subtracts the old LoRA weights and then adds the LoRA weights of the new task in the study by Hu et al. [14], we support multiple LoRA arms during inference at the same time, as shown in Figure 3. They share the same LLM memory as long as the compute resources allow.
@@ III-C Meta Planning
Instead of responding solely to cognitive heuristics, we have a meta-learning decision (tree) model to decide on the best inference planning paths for user prompts. Training inputs are user prompts, their end-to-end copilot responses, and various models for agent-specific tasks. The inference planning includes the combination and ordering of agents, e.g. only using Phi-3.5[2] or going through product recognition agent, database agents, etc. Slightly different from classical decision trees, we allow visiting the same nodes and paths several times by enabling retries during the inference stage. Then default paths, e.g., human customer service pop-up windows, and pre-calibrated messages, will be triggered.
Figure 4 shows the meta-learning decision (tree) model for our copilots. Each node is an AI language agent, and the orchestrator assigns the current user prompt to complete certain tasks. The planning decisions of the Meta-learner need to be actionable, e.g., proceeding with an exact agent with specific settings, such as English and U.S. The node can be whether a Microsoft product entity, such as Teams Essential, Microsoft 365 Business Premium, is recognized by the product recognition agent, and how many product entities are detected. AMO includes BERT, Phi-3.5 and their task-specific LoRA arms. The planning paths are deterministic during the interactions with customers after we learn the best actions from the training data. Therefore, reinforcement learning[26, 54, 55] was not the focus of our current study.
@@ IV-B Code Compliance Copilot
Our E-Commerce Copilot provides tailored responses to user prompt ”Do I have to use Outlook?” based on the web page information and shopping status.
@@ IV-C Experimental Results
We conduct experiments of classification tasks using our proposed agentic orchestrator, BERT, Fine-tuned Phi-3.5, ChatGPT-4o, BM 25 + Sentence Transformer[56] for a growing number of agents in Table 5. After we train agentic orchestrator, BERT, Phi-3.5 models using 9 out of 13 classes for the M365 E-Commerce Copilot data set and 5 out of 10 classes for the code compliance copilot data set, we inference them on the test data adding 5 new agents one by one. Our agentic orchestrator performs better than Phi-3.5, ChatGPT-4o models indicate that it is more reasonable to treat multi-label or hierarchical text classification as a multi-level rating learning-to-rank task than a generation task.
We show the superior end-to-end performance of ROUGE-L [57], BERTScore[58], and the F-1 measures of multiple tasks using our proposed AMO, Fine-tuned Phi-3.5, ChatGPT-4o, and AutoGen against the baseline of our corporate GPT-3.5 Turbo model in Table I.
@@
Extension for Agentic Orchestrator 0.96 0.85 0.82
@@ V Extension to Hierarchical Text Classification Results
To demonstrate why we chose uRank loss as the objective function instead of classification loss functions in our proposed agentic orchestrator, we combined uRank with an MPNet Sentence Transformer [59] and applied this combination to a hierarchical text classification [60] data set [^3] https://www.kaggle.com/datasets/kashnitsky/hierarchical-text-classification. This Amazon product review dataset contains 40k training records and 10k validation records. There are structured classes of three-levels: 6 “level 1” classes, 64 “level 2” classes, and 464 “level 3” classes, which inherit semantic hierarchies of relevance ratings. In table II, we report the superior results of our experimental setup compared to popular supervised text classifiers using transformers [61], sequence generation few-shot learners [^4] https://github.com/microsoft/autogen using Phi-4 model [^5] https://huggingface.co/microsoft/phi-4 and ChatGPT 4. For sequence generation of few-shots learners, we use the RAG method, which retrieves top-20 related records in training data using embeddings generated by the same Sentence Transformer for fair comparison. While text classifiers require three or more separate models for the three levels, and the sequence generation few-shot learners require multiple LLM calls by design, the levels of classes contribute to the data argumentation of our agentic orchestrator extension. It was interesting to find that, for text classification tasks, the uRank objective function and even the classical softmax-style text classifiers generally outperformed few-shot approaches using LLMs.
@@ VII Conclusion
We have proposed AMO that can orchestrate user prompts to the best agents and can resolve the uncertain top k selections for production autonomy. We have independent LoRA fine-tuning processes for various tasks, and the trained LoRA weights are used for those tasks while sharing the orignal base model during the inference stage. Finally, we demonstrate the efficiency of our meta-learning planner, selecting different inference strategies based on user prompts. We have shown the method is effective in Microsoft Copilot cases such as M365 E-Commerce Copilot and code compliance copilot. This architecture can also be extended to any multi-agent AI systems in various industries such as travel, online merchant, etc. We plan to explore reinforcement learning fine-tuning and choosing different arms [62, 63], reasoning and actions among collaborative intelligence[64, 65], federate learning at the copilot level [66].

### v2

@@ I Introduction
Microsoft copilots are designed to handle multiple work tasks by coordinating large language models (LLMs) [dubey2024llama3herdmodels, abdin2024phi3technicalreporthighly, mohamadi2023chatgptagegenerativeai, Dong_2022], Microsoft 365 productivity apps, and more. LLMs are impressive at handling general tasks, such as drafting documents, question answering, and summarization. Due to data compliance and boundaries, fine-tuning of customer data with web-based LLMs, such as ChatGPT, can be prohibited. Therefore, keeping the knowledge of an LLM up-to-date can be challenging.
@@
To provide a customized in-domain assistant for Microsoft customers, language agents can neatly extend the generalizability of LLMs in private domains [10.1145/3627673.3679105, yao2024tau, Wang_2024, zhou2023agentsopensourceframeworkautonomous] and connect with applications, knowledge graphs, databases, other copilots, etc. The explosion of adapting various agents in particular tasks or domains poses challenges for 1) an effective orchestrator that decomposes tasks from user prompts and assigns them to growing agents, 2) efficient deployment of the foundation model and agents, and 3) inference planning among agents, e.g., which agents should be used next, and what messages can be shared across agents.
Regarding the first challenge, we propose a trained multi-level rating learning-to-rank model (with graded relevance) to orchestrate the prompts to the right agents, which is generally taken as a hierarchical text classification problem [zhang2024teleclass, vaswani2023attentionneed, chen2019bert]. The agents can be “ask for price”, “compare products”, “contact human customer support”, etc. as in our M365 E-Commerce Copilot, and can be ”encoding character issue”, ”new character issue” as in the code compliance copilot. An orchestrator can be at the copilot level, e.g., it selects the right plugins  [^1] https://learn.microsoft.com/en-us/microsoft-365-copilot/extensibility/orchestrator with agents, and it can select the right models based on the business intents of user prompts. As the number of agents continues to grow, the use of a text classification model [liu2019robertarobustlyoptimizedbert, devlin-etal-2019-bert] can have difficulties in scalability, and the cosine similarity method can struggle when the descriptions of agents have overlaps or when user prompts are ambiguous.
Regarding the second challenge, we propose a memory-efficient inference framework, LoRA[hu2021loralowrankadaptationlarge] arms, where each arm handles one task. Hosting multiple LLMs requires substantial memory[vyas2024autonomousindustrialcontrolusing, zhang2024chainbuddyaiagentgenerating]. Inspired by the LoRA framework, we share the base LLM model among tasks and train LoRA weights for individual tasks.
Regarding the third challenge, we propose a novel meta-learning[lee-etal-2022-meta] decision tree model that trains the best inference model combinations and orderings for different prompts. The well-known successful framework, retrieval-augmented generation (RAG), is a simplified two-step inference planning based on human heuristics that can be taken as a web search agent and LLM question-answering agent. When a complex copilot service is composed of multiple agents and combinations of skills/models, how we effectively choose an inference strategy for individual user prompts requires a meta-learning approach. Current planning strategies, such as AutoGen[dibia2024autogen, porsdam2024autogen, wuautogen, zhu2023autogen, wu2023autogen], CrewAI, and TaskWeaver[barbarroxa2024benchmarking], rely on agentic workflows and task-specific prompts, which are generally based on cognitive architectures[sumers2023cognitive, ghosal2025doesthinkinghelpunderstanding, deng2025simura, kang2024self].
@@ II-A Task Distribution and Semantic Learning-to-rank
LLMs have been attempted in few-shot text classification tasks[li-etal-2024-pdameta, zhang-etal-2022-prompt-based] and few-shot learning-to-rank models[qin-etal-2024-large, jiang-etal-2023-llm, guo2022rankdnnlearningrankfewshot, sinhababu2024fewshotpromptingpairwiseranking]
Previous studies have shown that learning-to-rank can be used explicitly and implicitly in various language tasks [10.1145/3543873.3584621, karpukhin-etal-2020-dense]. However, they were limited to binary ratings. We apply semantic learning to rank to multi-level rating scenarios. We utilize the natural description of classification labels as the learning-to-rank “candidate documents”, different from the 0 to $n$ classes using one-hot encoding in softmax.
@@ II-B Multi-task Learning
Multi-task learners generally share a common module [zhang-etal-2023-survey, chen2019bert]. Previous studies, such as MT-DNN[bhattacharjee2022multendtoendmultitasklearning] and MulT[liu2019multi] have focused on training multiple tasks at the same time by sampling one task at a time using a neural network with shared layers and task-specific layers. This method can suffer performance drops in subsets of tasks, and it is not easy to efficiently update models based on individual task requirements. The authors of [stickland2019bert] proposed projected attention layers (PALs) in parallel to each layer for BERT [devlin-etal-2019-bert], which learns task-specific information in the additional parameters. The idea is similar to LoRA, which uses low-rank matrices to reduce the number of parameters for attention in task-specific layers. However, when we update the training data for one task, the performance of other tasks also changes, which is not good for updating the models, especially considering in many cases in industry we want to update the model for one task without affecting the performance for other tasks. We show that we achieve impressive individual and end-to-end task performance leveraging LoRA fine-tuning and inference memory optimization in Sections III-B - III-C and IV-C.
@@ II-C Agentic Workflow
Unlike chain of thoughts[wei2023chainofthoughtpromptingelicitsreasoning], tree of thoughts [yao2024tree], or tree search [light2024strategistlearningstrategicskills, yu2024exactteachingaiagents] that address LLM inference planning, our meta-learning planning strategy is a trained model of models/agents.
A well-known cognitive heuristics RAG [edge2024localglobalgraphrag] plans a knowledge retrieval then LLM generation. Although great reinforcement learning approaches, for example, React [yao2022react] and Reflection [shinn2023reflexionlanguageagentsverbal], and Self-critic [saunders2022selfcritiquingmodelsassistinghuman] have been attempted for complex systems, how we quickly update the planning when new agents are added remains an open problem. Our proposed AMO novelly solves this problem of continually adding new agents.
Another good cognitive heuristic is ensemble modeling. Each record is processed by the same models using bagging, stacking, etc. Our AMO, in contrast, teaches the methodology of training a model [ha-etal-2023-meta, 10.5555/3600270.3602999, chen-etal-2022-meta, zhang-etal-2022-prompt-based] and is Conversable; agents can share information with each other.
@@ III-A Agentic Orchestrator of Top-k Selections
Therefore, we design an orchestrator that can seamlessly take new agents on the basis of their agent/class labels. Figure 2 demonstrates that we convert agent/class descriptions to semantic embeddings offline and use them to train a listwise multi-level learning-to-rank model [10.1145/1963405.1963459, chen-etal-2020-hierarchical, 10.1145/3336191.3371814]. The model targets higher relatedness prediction scores for agents that are ranked higher than others. In particular, when there is only one correct item in a ranking list or classification group, such as ‘[0, 1, 0, 0]’, the gradients of ListMLE-style loss functions and Softmax are equivalent. However, the former is more applicable to multi-level ratings or hierarchical classes. The user prompt and the agents use the same embedding model, which is generic. We were able to achieve stable performance using BERT and Sentence Transformers as the foundation models, even more stable than using softmax classification trained for all classes.
Each batch contains the trainable embedding of $q$ a user prompt, fixed embeddings of $D^{q}=\{d\}$ the set of candidate agent descriptions associated with $q$, which can differ in batches, and $R(D^{q})=\{r(d)\}$ the relatedness rating for $D^{q}$, ranging over all real numbers (showing that the previous study in [10.1145/3336191.3371814] can be generalized). We maximize the likelihood of selecting a document or class $d$ from $c$ = $d\cup\tilde{s}$:
@@
The rating of every document in $\tilde{s}$ is less than r(d)r(d)[10.1145/3336191.3371814]. We finalize our general learning to rank loss function uRank loss by adding a weight factor of $2^{r(d)}-1$. We list and answer the following core questions accordingly.
@@
LoRA-Arms are trained independently during fine-tuning and can be used simultaneously during inference.
@@ III-B LoRA Arms
Slightly different from LoRA memory optimization that subtracts the old LoRA weights and then adds the LoRA weights of the new task in the study by Hu et al. [hu2021loralowrankadaptationlarge], we support multiple LoRA arms during inference at the same time, as shown in Figure 3. They share the same LLM memory as long as the compute resources allow.
@@ III-C Meta Planning
Instead of responding solely to cognitive heuristics, we have a meta-learning decision (tree) model to decide on the best inference planning paths for user prompts. Training inputs are user prompts, their end-to-end copilot responses, and various models for agent-specific tasks. The inference planning includes the combination and ordering of agents, e.g. only using Phi-3.5[abdin2024phi3technicalreporthighly] or going through product recognition agent, database agents, etc. Slightly different from classical decision trees, we allow visiting the same nodes and paths several times by enabling retries during the inference stage. Then default paths, e.g., human customer service pop-up windows, and pre-calibrated messages, will be triggered.
Figure 4 shows the meta-learning decision (tree) model for our copilots. Each node is an AI language agent, and the orchestrator assigns the current user prompt to complete certain tasks. The planning decisions of the Meta-learner need to be actionable, e.g., proceeding with an exact agent with specific settings, such as English and U.S. The node can be whether a Microsoft product entity, such as Teams Essential, Microsoft 365 Business Premium, is recognized by the product recognition agent, and how many product entities are detected. AMO includes BERT, Phi-3.5 and their task-specific LoRA arms. The planning paths are deterministic during the interactions with customers after we learn the best actions from the training data. Therefore, reinforcement learning[deng2025simura, jha2024neural, hou2504thinkprune] was not the focus of our current study.
@@ IV-B Code Compliance Copilot
Our E-Commerce Copilot provides tailored responses to user prompt “Do I have to use Outlook?” based on the web page information and shopping status.
@@ IV-C Experimental Results
We conduct experiments of classification tasks using our proposed agentic orchestrator, BERT, Fine-tuned Phi-3.5, ChatGPT-4o, BM 25 + Sentence Transformer[reimers2019sentencebertsentenceembeddingsusing] for a growing number of agents in Table 5. After we train agentic orchestrator, BERT, Phi-3.5 models using 9 out of 13 classes for the M365 E-Commerce Copilot data set and 5 out of 10 classes for the code compliance copilot data set, we inference them on the test data adding 5 new agents one by one. Our agentic orchestrator performs better than Phi-3.5, ChatGPT-4o models indicate that it is more reasonable to treat multi-label or hierarchical text classification as a multi-level rating learning-to-rank task than a generation task.
We show the superior end-to-end performance of ROUGE-L [lin-2004-rouge], BERTScore[zhang2020bertscoreevaluatingtextgeneration], and the F-1 measures of multiple tasks using our proposed AMO, Fine-tuned Phi-3.5, ChatGPT-4o, and AutoGen against the baseline of our corporate GPT-3.5 Turbo model in Table I.
@@
Extension for Agentic Orchestrator 0.96 0.85 0.83
@@ V Extension to Hierarchical Text Classification Results
To demonstrate why we chose uRank loss as the objective function instead of classification loss functions in our proposed agentic orchestrator, we combined uRank with an MPNet Sentence Transformer [song2020mpnet] and applied this combination to a hierarchical text classification [aly2019hierarchical] data set [^3] https://www.kaggle.com/datasets/kashnitsky/hierarchical-text-classification. This Amazon product review dataset contains 40k training records and 10k validation records. There are structured classes of three-levels: 6 “level 1” classes, 64 “level 2” classes, and 464 “level 3” classes, which inherit semantic hierarchies of relevance ratings. In table II, we report the superior results of our experimental setup compared to popular supervised text classifiers using transformers [meng2019weakly], sequence generation few-shot learners [^4] https://github.com/microsoft/autogen using Phi-4 model [^5] https://huggingface.co/microsoft/phi-4 and ChatGPT 4. For sequence generation of few-shots learners, we use the RAG method, which retrieves top-20 related records in training data using embeddings generated by the same Sentence Transformer for fair comparison. While text classifiers require three or more separate models for the three levels, and the sequence generation few-shot learners require multiple LLM calls by design, the levels of classes contribute to the data argumentation of our agentic orchestrator extension. It was interesting to find that, for text classification tasks, the uRank objective function and even the classical softmax-style text classifiers generally outperformed few-shot approaches using LLMs.
@@ VII Conclusion
We have proposed AMO that can orchestrate user prompts to the best agents and can resolve the uncertain top k selections for production autonomy. We have independent LoRA fine-tuning processes for various tasks, and the trained LoRA weights are used for those tasks while sharing the orignal base model during the inference stage. Finally, we demonstrate the efficiency of our meta-learning planner, selecting different inference strategies based on user prompts. We have shown the method is effective in Microsoft Copilot cases such as M365 E-Commerce Copilot and code compliance copilot. This architecture can also be extended to any multi-agent AI systems in various industries such as travel, online merchant, etc. We plan to explore reinforcement learning fine-tuning and choosing different arms [cruz2023reinforcementlearningfinetuninglanguage, 10.1145/3546790.3546804], reasoning and actions among collaborative intelligence[10.5555/3635637.3663215, zhang2024raftadaptinglanguagemodel], federate learning at the copilot level [10.1145/3664650].